# Interferometry of non-Abelian band singularities and Euler class topology

### v1

@@ 00 $\pi$ $\cos^{2}{\phi}$ $\sin^{2}{\phi}$ same
Figure 3: a) Illustration of the consecutive deflection method. Along the path joining the nodes, a relative dynamic phase $\alpha^{\prime}$ is developed between the bands. b) Starting from band $n=1$, the final population $p_{1}$ in this band oscillates as a function of $\theta_{1}$, which shifts right (left) for increasing $\theta_{2}$ values if the frame charges are same (opposite), here demonstrated in Kagome lattice for $(K,K^{\prime})$ nodes (2). c) The phase shift direction is deduced from Eqs.(5,6), after determining $\cos{\alpha^{\prime}}<0$ at double acceleration with $\theta_{2}=\pi/2$.
@@
Method 2: Consecutive Deflection.— Our second method does not require an interferometry loop and offers the advantage of not increasing in complexity when symmetries are removed. This relies on a single wavepacket entering and exiting the two targeted nodes between bands $n$ and $n+1$ consecutively as depicted in Fig. 3a. Starting in band $n$, the final population $p_{n}$ in this band depends on the frame-rotation angles $\phi_{1}$ and $\phi_{2}$, as well as the relative dynamic phase $\alpha^{\prime}$ developed in between the nodes, which a priori complicates deducing the frame charges  [^5] Any relative dynamic phases appearing along the pre/post-node sections are irrelevant since they will not affect the populations.. We circumvent this by expressing it after some algebra [44] as $p_{n}=\left[A\cos{\left(2\phi_{1}-\beta\right)}+1\right]/2$, where
$A^{2}=\cos^{2}{\left(2\phi_{2}\right)}+\cos^{2}{\left(\alpha^{\prime}\right)}\sin^{2}{\left(2\phi_{2}\right)},$ (5)
$\tan{\beta}=-\cos{\left(\alpha^{\prime}\right)}\tan(2\phi_{2}).$ (6)
@@
Namely, the band population varies with $\phi_{1}$ sinusoidally with an amplitude $A$ and phase $\beta$. Crucially, the phase shift in Eq. (6) depends on whether the frame winds in the same/opposite direction around the second node $(\phi_{2})$, corresponding to relative same/opposite frame charges, provided that $\cos{\left(\alpha^{\prime}\right)}$ is known. The latter can be extracted from the amplitude (5), which reveals only $\cos^{2}(\alpha^{\prime})$, by performing the experiment at double acceleration to halve the dynamic phase. Obtaining $\cos^{2}{\left(\alpha^{\prime}/2\right)}$ thus yields $\cos{\alpha^{\prime}}=2\cos^{2}{\left(\alpha^{\prime}/2\right)}-1$. Alternatively, the dynamic phase can be also directly calculated from energy measurements.
@@
We numerically demonstrate this method by using the $(K,K^{\prime})$ nodes in Fig. 3 and other nodes in more complicated settings in SM [44], where $p_{n=1}$ oscillates as a function of the first turning angle $\theta_{1}=2\phi_{1}$. Repeating the consecutive deflection for different turning angles in node 2 induces a phase shift, direction of which reveals the relative frame charges targeted. Here, for $\cos\alpha^{\prime}<0$ extracted from the amplitude (5), oscillations shift to right (left) for similar (opposite) frame charges  [^6] For $\cos\alpha^{\prime}>0$, these shifts are reversed.
@@
Discussion.— We have presented two interferometric schemes to extract non-Abelian frame charges of band degeneracies of real-valued Hamiltonians, which require control on lattice acceleration and path precision that have been already demonstrated in state-of-the-art ultracold atom experiments [34, 32]. While the first method involves only two measurements –the actual interferometry and in general a reference loop if dynamic phases are unknown [44] which is not necessary in the presence of lattice symmetries–, the second method relies on several measurements for a range of $\theta_{1,2}$ values to detect the phase shift but applies independently of underlying lattice symmetries. We note that pseudospins in Method 1 can be employed for guiding atoms along an interferometry loop [32]. This can be further combined with band specific readout of atoms to extract individual terms in Table 1. Method 2 can be fine-tuned for a specific experiment as well. For example, if the rate of frame winding around individual nodes is known (e.g. uniform under $C_{6}$ symmetry [44]), one requires only three measurements: Two for observing $p_{n}$ under double acceleration by choosing $|\phi_{1}|=|\phi_{2}|=\pi/4$ and reversing the sign of $\phi_{1}$, which yield $\sin^{2}{(\alpha^{\prime}/4)}$ and $1-\sin^{2}{(\alpha^{\prime}/4)}$ to find $\cos^{2}{(\alpha^{\prime}/2)}$. Repeating the first measurement at normal acceleration then discloses $1-\cos^{2}{(\alpha^{\prime}/2)}~{}\,(\cos^{2}{(\alpha^{\prime}/2)})$ for similar (opposite) charges.
@@
The interferometry and consecutive deflection methods reveal the relative frame charge along the path joining the nodes. Hence, by varying the route connecting the nodes, our techniques pave the way for directly demonstrating the non-Abelian and path dependent nature of braiding in experiment for the first time. Given the essential role of such charges in constituting Euler class and multi-gap topology [18, 17, 19], these results serve as a crucial route towards experimentally probing multi-gap topologies in and out of equilibrium [29].
@@ Acknowledgements.
Acknowledgments— O.B. thanks James Walkling and Calvin Hooper for productive discussions. We acknowledge discussions with Dan Stamper-Kurn and Adrien Bouhon. O.B. acknowledges funding from the Clarendon Fund, Merton College, and a Leverhulme Studentship. R.-J. S. acknowledges funding from a New Investigator Award, EPSRC grant EP/W00187X/1, a EPSRC ERC underwrite grant EP/X025829/1, and a Royal Society exchange grant IES/R1/221060 as well as Trinity College, Cambridge. F.N.Ü. acknowledges funding from the Marie Skłodowska-Curie programme of the European Commission Grant No 893915, Trinity College Cambridge, and thanks the Aspen Center for Physics for their hospitality, where this work was partially funded by a grant from the Sloan Foundation.
@@
In $\mathcal{C}_{2}\mathcal{T}$ symmetric systems, the eigenvectors are real and orthogonal, and thus at each quasimomentum form an orthonormal frame $\{\mathinner{|{u_{n}(k)}\rangle}\}_{n=1}^{N}$ [17, 22]. However, each eigenvector is only defined up to an overall sign $\pm$ (a gauge degree of freedom). Considering a closed path, since the initial and final Hamiltonians will be identical, the accumlated angle defines a ‘frame charge’ associated with the path. In particular, a frame charge may be nontrivial when nodes are enclosed. For example, for a path enclosing a single linear node between two bands (the ‘principal bands’), the eigenvectors corresponding to the principal bands undergo a $\pi$ rotation as we circle the node [17]. Such a $\pi$ winding means that it is not possible to define a globally smooth gauge and there is a line across which the gauge switches so the sign of these two bands changes. This line, similar to a branch cut, is called a Dirac string [18], and is demonstrated in Fig. 1. A Dirac string may either terminate on a node (joining a pair of nodes) or may cross the entire Brillouin zone (e.g. as in the anomalous Floquet phases introduced in Ref. [29]) after nodes are recombined over the Brillouin zone edge.
@@
The sign of (the quaternion charge of) the node refers to the direction in which the eigenvectors wind. While these signs can be changed through a gauge transformation, relative charge in a band subspace can be defined by smoothly fixing the gauge between the nodes. For two similarly-charged nodes in a given gap, the net frame winding around a path encircling them is $2\pi$, indicating an obstruction to the ability of these nodes to annihilate. For oppositely charged nodes, the net winding is 00, and the nodes may annihilate. The relative frame charges can be path dependent as demonstrated in Fig. S4, in which two nodes between the principal bands ($\mathinner{|{u_{1}}\rangle},\mathinner{|{u_{2}}\rangle}$) are labelled by ±$\pm i$, and one ‘adjacent’ node by $+k$. Suppose that within the enclosed region in (a), the two $i$ nodes have opposite charge. We now ask what the relative charge would be for the region on the other side of the $k$ node, as in (c). In fixing the gauge to be smooth within that region, we must move the red Dirac string outside of this region. Pulling this string through the $-i$ node reverses the signs of $\mathinner{|{u_{2}}\rangle},\mathinner{|{u_{3}}\rangle}$, and thus the winding of the $\mathinner{|{u_{1}}\rangle},\mathinner{|{u_{2}}\rangle}$ frame changes. As a result, within this region in (c) the $i$ nodes have the same charge. We stress that the relative stability is a gauge invariant physical quantity and even motivates a topological invariant, the patch Euler class. Indeed, considering region $\Gamma_{a}$ in Fig. S4c the $i$ nodes still can annihilate in this patch as one transverses the Dirac string of the adjacent gap, a similar analysis holds for considering region $\Gamma_{c}$ in Fig. S4a. This motivates a puzzle approach [39, 49] in which one fixes the gauge at a patch by referring to relative stability, e.g. on region $\Gamma_{a}$ and build up the global topological configuration using that the frame charge changes sign when crossing the Dirac strings residing in adjacent gaps as well as combination rules for Dirac strings [29].
@@
Figure S4: Demonstration of how the relative charge of two nodes depends on the region considered. Nodes and Dirac strings between bands 1 and 2 are in blue and labelled by $i$, those between bands 2 and 3 are in red and labelled by $k$. a) In the region $\Gamma_{a}$ the charges have opposite charge, with a patch Euler class $\chi=0$, or equivalently the winding of the frame along $\Gamma_{a}$ is 00. b) Pulling the Dirac strings through the node in the other gap change their charge, resulting in a smoothly defined gauge within $\Gamma_{c}$. c) In the region $\Gamma_{c}$ the nodes have the same charge and thus $\chi=1$. The frame winds by $2\pi$ along $\Gamma_{c}$ as a manifestation of the non-Abelian nature of the band nodes. We emphasise that each subfigure is the same physical situation, but each with a different gauge choice.
@@
The above labelling of the nodes with quaternion charges $i,j,k$ is not unique. Consider nodes only in a 3-band subspace, and suppose we circle a node between bands 1-2 and another between bands 2-3. Composing a $\pi$ rotation of $\mathinner{|{u_{1}}\rangle},\mathinner{|{u_{2}}\rangle}$ followed by a $\pi$ rotation of $\mathinner{|{u_{2}}\rangle},\mathinner{|{u_{3}}\rangle}$, is in fact equivalent to a $\pi$ rotation of $\mathinner{|{u_{3}}\rangle},\mathinner{|{u_{1}}\rangle}$ [29]. Considering different pairs of bands to get equivalent relations, and noting that a trivial or $2\pi$ winding in any gap is equivalent, we can see that the set of charges can be described by the quaternion group with charge composition described by quaternion multiplication rules. More explicitly, we define the charges of nodes e.g. of bands 1-2 as ±$\pm i$ (sign depending on direction of winding), bands 2-3 as ±$\pm k$, bands 3-1 as ±$\pm j$, the trivial winding as 1 and the $2\pi$ winding as -1. Then charge composition is reflected by
@@
.2 II. Details for Method 1: Interferometry

### v2

@@ 00 $\pi$ $\cos^{2}{\phi^{\mathrm{frame}}}$ $\sin^{2}{\phi^{\mathrm{frame}}}$ same
Figure 3: a) Illustration of the consecutive deflection method. Between the nodes, a relative dynamic phase $\alpha^{\prime}$ is developed. b) Starting from band $n=1$, the final population $p_{1}$ oscillates as a function of $\phi^{\mathrm{path}}_{1}$. The phase-shift direction for increasing $\phi^{\mathrm{path}}_{2}$ depends on the relative frame charge and sign of $\cos{\alpha^{\prime}}$, here demonstrated in Kagome lattice for $(K,K^{\prime})$ nodes (2). c) $\cos{\alpha^{\prime}}<0$ is deduced from Eq. (5) at double acceleration with $\phi^{\mathrm{frame}}_{2}=\pi/4$.
@@
Method 2: Consecutive Deflection.— Our second method does not require an interferometry loop and offers the advantage of not increasing in complexity when symmetries are removed. This relies on a single wavepacket entering and exiting the two targeted nodes [in the same way as in Eq. (Interferometry of non-Abelian band singularities and Euler class topology)] between bands $(n,\,n+1)$ consecutively [Fig. 3a]. Starting in band $n$, the final population $p_{n}$ in this band depends on the frame-rotation angles $(\phi^{\mathrm{frame}}_{1},\phi^{\mathrm{frame}}_{2})$, as well as the relative dynamic phase $\alpha^{\prime}$ developed in between the nodes, which a priori complicates deducing the frame charges  [^6] Any relative dynamic phase appearing along the pre/post-node sections are irrelevant since they will not affect the populations.. We circumvent this by expressing it after some algebra [42] as $p_{n}\!\!=\!\!\left[A\cos{\left(2\phi^{\mathrm{frame}}_{1}\!-\!\beta\right)}\!+\!1\right]\!/2$, where
$A^{2}=\cos^{2}{\left(2\phi^{\mathrm{frame}}_{2}\right)}+\cos^{2}{\left(\alpha^{\prime}\right)}\sin^{2}{\left(2\phi^{\mathrm{frame}}_{2}\right)},$ (5)
$\tan{\beta}=-\cos{\left(\alpha^{\prime}\right)}\tan(2\phi^{\mathrm{frame}}_{2}).$ (6)
@@
The band population varies with $\phi^{\mathrm{frame}}_{1}$ sinusoidally with an amplitude $A$ and phase $\beta$. Crucially, the phase shift in Eq. (6) depends on whether the frame winds in the same/opposite direction around the second node $(\phi^{\mathrm{frame}}_{2})$, corresponding to same/opposite frame charges, provided that $\cos{\left(\alpha^{\prime}\right)}$ is known. The latter can be extracted from the amplitude (5), which reveals only $\cos^{2}(\alpha^{\prime})$, by performing the experiment at double the acceleration to halve the dynamic phase. Obtaining $\cos^{2}{\left(\alpha^{\prime}/2\right)}$ thus yields $\cos{\alpha^{\prime}}=2\cos^{2}{\left(\alpha^{\prime}/2\right)}-1$. Alternatively, the dynamic phase can be directly calculated from energy measurements [46, 31].
@@
We numerically demonstrate this method by using the $(K,K^{\prime})$ nodes in Fig. 3 (see SM [42] for more complicated settings), where $p_{n=1}$ oscillates as a function of the first turning angle $\phi^{\mathrm{path}}_{1}=2\phi^{\mathrm{frame}}_{1}$. Repeating for different turning angles in node 2 induces a phase shift, direction of which reveals the relative frame charges. Fig. 3c reveals $\cos\alpha^{\prime}<0$ in both cases, and therefore the phase $(\beta)$ increases (decreases) for similar (opposite) frame charges.
@@
Discussion.— Our two interferometric schemes for extracting non-Abelian frame charges of real Hamiltonians require control on lattice acceleration and path precision that have been already demonstrated in ultracold atom experiments [34, 32]. While the first method involves two measurements–the actual interferometry and in general a reference loop if dynamic phases differ along each path [42]–, the second method relies on several measurements for a range of $\phi^{\mathrm{path}}_{1,2}$ values to detect the phase shift but applies independently of underlying lattice symmetries. We note that pseudospins in Method 1 can be employed for guiding atoms along an interferometry loop [32] and further combining with band specific readouts of atoms reveals individual terms in Table 1. Method 2 can be fine-tuned for a specific experiment as well. For example, for known frame windings around individual nodes (e.g. uniform under $C_{6}$ symmetry [42]), one requires only three measurements: Two for observing $p_{n}$ under double acceleration by choosing $|\phi^{\mathrm{frame}}_{1}|=|\phi^{\mathrm{frame}}_{2}|=\pi/4$ and reversing the sign of $\phi^{\mathrm{frame}}_{1}$, which yield $\sin^{2}{(\alpha^{\prime}/4)}$ and $1-\sin^{2}{(\alpha^{\prime}/4)}$ to find $\cos^{2}{(\alpha^{\prime}/2)}$. Repeating the first measurement at normal acceleration then discloses $1-\cos^{2}{(\alpha^{\prime}/2)}~{}\,(\cos^{2}{(\alpha^{\prime}/2)})$ for similar (opposite) charges. We further demonstrate our techniques’ robustness under experimental imperfections in [42], putting them within reach of state-of-the-art experiments [32, 34].
@@
The interferometry and consecutive deflection methods reveal the relative frame charge along the path joining the nodes. Hence, by varying the route connecting the nodes, our techniques pave the way for directly demonstrating the non-Abelian and path dependent nature of braiding in experiments. Given the essential role of such charges in constituting Euler class [18, 17, 19], these serve as a crucial route towards experimentally probing multi-gap topologies in and out of equilibrium [29].
@@ Acknowledgements.
Acknowledgments— O.B. thanks James Walkling and Calvin Hooper for productive discussions. We acknowledge discussions with Dan Stamper-Kurn and Adrien Bouhon. O.B. acknowledges funding from the Clarendon Fund, Merton College, and a Leverhulme Studentship. R.-J. S. acknowledges funding from a New Investigator Award, EPSRC grant EP/W00187X/1, a EPSRC ERC underwrite grant EP/X025829/1, and a Royal Society exchange grant IES/R1/221060 as well as Trinity College, Cambridge. F.N.Ü. acknowledges funding from the Marie Skłodowska-Curie programme of the European Commission Grant No 893915, Simons Investigator Award [Grant No. 511029], Trinity College Cambridge, and thanks the Aspen Center for Physics for their hospitality, where this work was partially funded by a grant from the Sloan Foundation.
@@
Frame charges, Dirac strings, and their non-Abelian properties are described in-depth in Ref. [17]. Here we provide a brief exposition to the key concepts.
@@
In $\mathcal{C}_{2}\mathcal{T}$ symmetric systems, the eigenvectors are real and orthogonal, and thus at each quasimomentum form an orthonormal frame $\{\mathinner{|{u_{n}(k)}\rangle}\}_{n=1}^{N}$ [17, 22]. However, each eigenvector is only defined up to an overall sign $\pm$ (a gauge degree of freedom). Considering a closed path, since the initial and final Hamiltonians will be identical, the accumlated angle defines a ‘frame charge’ associated with the path. In particular, a frame charge may be nontrivial when nodes are enclosed. For example, for a path enclosing a single linear node between two bands (the ‘principal bands’), the eigenvectors corresponding to the principal bands undergo a $\pi$ rotation as we circle the node [17]. Such a $\pi$ winding means that it is not possible to define a globally smooth gauge and there is a line across which the gauge switches so the sign of these two bands changes. This line, similar to a branch cut, is called a Dirac string [18], and is demonstrated in Fig. 1(b). A Dirac string may either terminate on a node (joining a pair of nodes) or may cross the entire Brillouin zone (e.g. as in the anomalous Floquet phases introduced in Ref. [29]) after nodes are recombined over the Brillouin zone edge.
We can see this winding more explicitly as follows. For a linear node, expanding in wavevectors about the degenerate point in the subspace corresponding to these bands, the Hamiltonian must take the form [17]
$H_{2}=\sum_{i=1}^{2}\sum_{j=0}^{3}k_{i}h_{ij}\sigma_{j},$ (S1)
where the coefficients $h_{ij}$ are real and $h_{i2}=0$. Neglecting the identity component (which does not affect the eigenstates),
$H_{2}=\begin{pmatrix}k_{i}h_{i3}&k_{i}h_{i1}\\ k_{i}h_{i1}&-k_{i}h_{i3}\end{pmatrix}.$ (S2)
Making the linear change of basis
$\begin{pmatrix}\kappa_{1}\\ \kappa_{2}\end{pmatrix}=\begin{pmatrix}h_{13}&h_{23}\\ h_{11}&h_{21}\end{pmatrix}\begin{pmatrix}k_{1}\\ k_{2}\end{pmatrix},$ (S3)
we finally have
$H_{2}=\begin{pmatrix}\kappa_{1}&\kappa_{2}\\ \kappa_{2}&-\kappa_{1}\end{pmatrix}=|\kappa|\left[\cos{\theta}\sigma_{1}+\sin{\theta}\sigma_{3}\right],$ (S4)
The eigenstates take the form
$\mathinner{|{u_{1}}\rangle}=\begin{pmatrix}\cos{\theta/2}\\ -\sin{\theta/2}\end{pmatrix},\mathinner{|{u_{2}}\rangle}=\begin{pmatrix}\sin{\theta/2}\\ \cos{\theta/2},\end{pmatrix}$ (S5)
i.e. the frame angle is $\phi^{\mathrm{frame}}=\theta/2$. As we move around the node $\theta$ varies from 00 to $2\pi$, and therefore the frames wind by an angle of $\pi$. This winding is demonstrated in Fig. 1b in the main text. The change of basis in Eqn. (S3) is not necessarily orthogonal, and thus a uniform winding of $\theta$ here does not always correspond to a uniform winding of the frames in terms of the angle in the original $(k_{1},k_{2})$ basis.
@@
The sign of (the quaternion charge of) the node refers to the direction in which the eigenvectors wind. While the sign of a given node can be changed through a gauge transformation, relative charge in a band subspace can be defined by smoothly fixing the gauge between the nodes. For two similarly-charged nodes in a given gap, the net frame winding around a path encircling them is $2\pi$, indicating an obstruction to the ability of these nodes to annihilate. For oppositely charged nodes, the net winding is 00, and therefore the nodes may annihilate. The relative frame charges can be path dependent as demonstrated in Fig. S4, in which two nodes between the principal bands ($\mathinner{|{u_{1}}\rangle},\mathinner{|{u_{2}}\rangle}$) are labelled by ±$\pm i$, and one ‘adjacent’ node by $+k$. Suppose that within the enclosed region in (a), the two $i$ nodes have opposite charge. We now ask what the relative charge would be for the region on the other side of the $k$ node, as in (c). In fixing the gauge to be smooth within that region, we must move the red Dirac string outside of this region. Pulling this string through the $-i$ node reverses the signs of $\mathinner{|{u_{2}}\rangle},\mathinner{|{u_{3}}\rangle}$, and thus the winding of the $\mathinner{|{u_{1}}\rangle},\mathinner{|{u_{2}}\rangle}$ frame changes. As a result, within this region in (c) the $i$ nodes have the same charge. We stress that the relative stability is a gauge invariant physical quantity and even motivates a topological invariant, the patch Euler class. Indeed, considering region $\Gamma_{a}$ in Fig. S4c the $i$ nodes still can annihilate in this patch as one transverses the Dirac string of the adjacent gap, a similar analysis holds for considering region $\Gamma_{c}$ in Fig. S4a. This motivates a puzzle approach [40, 49] in which one fixes the gauge at a patch by referring to relative stability, e.g. on region $\Gamma_{a}$ and build up the global topological configuration using that the frame charge changes sign when crossing the Dirac strings residing in adjacent gaps as well as combination rules for Dirac strings [29].
@@
Figure S4: Demonstration of how the relative charge of two nodes depends on the region considered. Nodes and Dirac strings between bands 1 and 2 are in blue and labelled by $i$, those between bands 2 and 3 are in red and labelled by $k$. a) In the region $\Gamma_{a}$ the nodes have opposite charge, with a patch Euler class $\chi=0$, or equivalently the winding of the frame along $\Gamma_{a}$ is 00. b) Pulling the Dirac strings through the node in the other gap change their charge, resulting in a smoothly defined gauge within $\Gamma_{c}$. c) In the region $\Gamma_{c}$ the nodes have the same charge and thus $\chi=1$. The frame winds by $2\pi$ along $\Gamma_{c}$ as a manifestation of the non-Abelian nature of the band nodes. We emphasise that each subfigure is the same physical situation, but each with a different gauge choice.
@@
The above labelling of the nodes with quaternion charges $i,j,k$ is not unique. Consider nodes only in a 3-band subspace, and suppose we circle a node between bands 1-2 and another between bands 2-3. Composing a $\pi$ rotation of $\mathinner{|{u_{1}}\rangle},\mathinner{|{u_{2}}\rangle}$ followed by a $\pi$ rotation of $\mathinner{|{u_{2}}\rangle},\mathinner{|{u_{3}}\rangle}$, is in fact equivalent to a $\pi$ rotation of $\mathinner{|{u_{3}}\rangle},\mathinner{|{u_{1}}\rangle}$ [29]. Considering different pairs of bands to get equivalent relations, and noting that a trivial or $2\pi$ winding in any gap is equivalent, we can see that the set of charges can be described by the quaternion group with charge composition described by quaternion multiplication rules. More explicitly, we define the charges of nodes e.g. of bands 1-2 as ±$\pm i$ (sign reflecting the opposite direction of winding), bands 2-3 as ±$\pm k$, bands 3-1 as ±$\pm j$, the trivial winding as 1 and the $2\pi$ winding as -1. Then charge composition is reflected by
@@
.2 II. Motion through the Brillouin Zone
Figure S5: A schematic demonstrating the occupation of eigenstates as we enter and exit a linear node at $\phi^{\mathrm{frame}}=0$, $\pi$, and some intermediate angle. On the left we repeat the demonstration of frame winding from the main text; the eigenstates corresponding to the node rotate by $\pi$ as we circle the node, resulting in the final band populations indicated on and below the cones.
We emphasise that while moving the wavepacket through the BZ and for the resulting superposition given in Eq.(3) of the main text, the motion will be adiabatic for small enough acceleration and away from any nodes. This adiabaticity with respect to the third band, or all other bands outside of the Euler subspace can be in general easily satisfied as these bands will be energetically separated. We note that also within the Euler subspace, no excitations occur until the band touching point due to the orthonormality of the frame eigenstates along the in-going path and the wavefunction will acquire only a dynamic phase as will be shown below. Since the eigenstates are real, there will be no Berry phase acquired along these straight paths.
In this section we now give a detailed discussion of the dynamics of a wavepacket under an applied force, before applying this to the motion in and out of a node. We discuss the roles of adiabaticity and the effect of ‘missing’ the node by a small amount. Several of these concepts are also discussed in the Supplementary Material for Ref. [34].
.2.1 Motion under an applied force
When an external force $\bm{F}$ is applied to particle which otherwise has Hamiltonian $H$, the Schrödinger equation is
$i\partial_{t}\mathinner{|{\psi}\rangle}=(H-\bm{F}\cdot\bm{r})\mathinner{|{\psi}\rangle}.$ (S8)
$\mathinner{|{\psi}\rangle}=e^{i\bm{k}\cdot\bm{r}}\mathinner{|{u(\bm{k})}\rangle},$ (S9)
where $\mathinner{|{u(\bm{k})}\rangle}$ is a superposition of bands
$\mathinner{|{u_{k}}\rangle}=\sum_{n}c_{n}\mathinner{|{u_{n}(\bm{k})}\rangle}.$ (S10)
The effect of the force is that $\dot{\bm{k}}=\bm{F}$ [50], such that through an applied force we can choose the path through momentum-space $\bm{k}(t)$. Therefore, the wavefunction will evolve as
$\mathinner{|{\psi(t)}\rangle}=e^{i\bm{k}(t)\cdot\bm{r}}\mathinner{|{u(t)}\rangle},$ (S11)
with
$\mathinner{|{u(t)}\rangle}=\sum_{n}c_{n}(t)\mathinner{|{u_{n}(\bm{k}(t))}\rangle}.$ (S12)
$i\partial_{t}\mathinner{|{u(t)}\rangle}=H_{\bm{k}(t)}\mathinner{|{u(t)}\rangle},$ (S13)
where
$H_{\bm{k}}=e^{-i\bm{k}\cdot\bm{r}}He^{i\bm{k}\cdot\bm{r}}$ (S14)
is the Bloch Hamiltonian, and $\bm{k}(t)$ now acts as a varying external parameter yielding a time-dependent Hamiltonian $H_{k(t)}$.
[…]
Since the bands are real, no Berry phase is acquired. When a node is encountered and the motion is no longer adiabatic, transitions between the bands can occur; this is explored in the next section.
.2.2 Motion through a node
Consider a single Dirac cone of infinite size and uniform winding, with Hamiltonian $H=v(k_{x}\sigma_{x}+k_{y}\sigma_{z})$. Around the node, the eigenstates take the form
$\mathinner{|{u_{1}(\phi^{\mathrm{path}})}\rangle}=\begin{pmatrix}\sin{\phi^{\mathrm{path}}/2}\\ -\cos{\phi^{\mathrm{path}}/2}\end{pmatrix},~{}\mathinner{|{u_{2}(\phi^{\mathrm{path}})}\rangle}=\begin{pmatrix}\cos{\phi^{\mathrm{path}}/2}\\ \sin{\phi^{\mathrm{path}}/2}\end{pmatrix}.$ (S16)
As we wind round the node once, taking $\phi^{\mathrm{path}}$ from $0\to 2\pi$, the frame angle $\phi^{\mathrm{frame}}=\phi^{\mathrm{path}}/2$ winds from $0\to\pi$.
Suppose that we take a path which enters and leaves the node at an angle $\phi^{\mathrm{path}}=2\phi^{\mathrm{frame}}$, corresponding to
$\boldsymbol{k}(t)=\begin{pmatrix}|\beta t|\cos{\phi^{\mathrm{frame}}}\\ \beta t\sin{\phi^{\mathrm{frame}}}\end{pmatrix},$ (S17)
where we have expressed this in terms of $\phi^{\mathrm{frame}}$ such that our following discussion also applies to nodes with non-uniform winding.
For $t<0$ and separately for $t>0$, the eigenstates are constant. As a result, independently of the acceleration $\beta$, excitations only occur at the node $t=0$. Starting in band $\mathinner{|{u_{1}}\rangle}$ at time $t_{0}<0$, the state evolves as
$\mathinner{|{u(t)}\rangle}=\begin{cases}e^{iv(t^{2}-t_{0}^{2})/2}\mathinner{|{u_{1}(-\phi^{\mathrm{frame}})}\rangle}&t<0\\ e^{ivt^{2}/2}\cos{\phi^{\mathrm{frame}}}\mathinner{|{u_{1}(\phi^{\mathrm{frame}})}\rangle}+e^{-ivt^{2}/2}\sin{\phi^{\mathrm{frame}}}\mathinner{|{u_{2}(\phi^{\mathrm{frame}})}\rangle}&t>0.\end{cases}$ (S18)
To understand this, notice that for $t<0$ we stay in an unchanging eigenstate $\mathinner{|{u_{1}(-\phi^{\mathrm{frame}})}\rangle}$ of the Hamiltonian, so time evolution only acquires a phase. For $t>0$, the eigenstates are now $\mathinner{|{u_{1/2}(+\phi^{\mathrm{frame}})}\rangle}$, and therefore we decompose our $t=0$ state in terms of the new eigenstates $\mathinner{|{u_{1}(-\phi^{\mathrm{frame}})}\rangle}=\cos{\phi^{\mathrm{frame}}}\mathinner{|{u_{1}(\phi^{\mathrm{frame}})}\rangle}-\sin{\phi^{\mathrm{frame}}}\mathinner{|{u_{2}(\phi^{\mathrm{frame}})}\rangle}$, before applying their respective dynamic phase factors.
Before moving on, let’s note some important points. First, since the bands eigenstates are always real, the $\pi$ Berry phase upon circling around the node and the $\pi$ Berry flux of the node is captured by the Dirac string, with our eigenstates picking up a minus sign as we wind once. Secondly, the acceleration $\beta$ here was irrelevant, due to the scale-invariance of the Dirac Hamiltonian and path. A more realistic picture will be discussed in the following section.
.2.3 Non-adiabatic motion through a node
In this section we consider what happens when our path ‘misses’ the node by a small amount, and demonstrate that for sufficiently small misses the results of the previous section still hold. We shall continue to study the same single Dirac cone as in the previous section, but modify the path slightly to
$\boldsymbol{k}(t)=\begin{pmatrix}k_{x}\\ k_{y}\end{pmatrix}=\begin{pmatrix}\sqrt{(\beta t\cos{\phi^{\mathrm{frame}}})^{2}+\delta^{2}}\\ \beta t\sin{\phi^{\mathrm{frame}}}\end{pmatrix}$ (S19)
where we now miss the node by an amount $\delta$ as illustrated in Fig. S6.
Figure S6: Population which remaining in the initial band after missing the node by some amount $\delta$, with an acceleration $\beta$. As long as $\delta$ is sufficiently small ($\delta^{2}/\beta\ll 1$), the populations and dynamic phases lie close to the ideal case.
Provided that $v\frac{\delta^{2}}{\beta}\ll 1$, the populations (and dynamic phases - since the difference in dynamic phase will be roughly the change in energy $v\delta$ over a time period $\frac{\delta}{\beta}$) shall be close to those for $\delta=0$, adiabatically. This is demonstrated in Fig. S6, in which the final population in the initial band is plotted as a function of angle $\phi^{\mathrm{frame}}$ for a range of $\delta$.
$\beta$ is itself bounded by the constraint that motion is fully adiabatic away from the node. This depends on the exact band structure in question, through the energy gaps between the nodes, the rate at which eigenvectors change, and the extent of the node. The extent of the node is simply the region over which the linear expansion in wavevector about the band touching point is valid, such that the rate of winding is fixed. Numerically we find that, for all nodes studied, the extent of the node is at least $\delta k\sim 0.1$. Intuitively, we require perfectly adiabatic motion outside of this region, but inside this circle we are trying to ‘project’ the wavefunction into the bands at some angle around the node, and therefore within this region non-adiabaticity is allowed (and indeed required for non-zero $\delta$ through $v\delta^{2}/\beta\ll 1$).
Further constraints arise from practical considerations. First, the wavepacket possesses finite spread in momentum space. For our methods to work, we require that the phase of the wavefunction is approximately the same over the entire wavepacket. The range of dynamic phases can be estimated from the range of average energies multiplied by the total time taken for the interferometry. In general this condition will be similar to the above condition, as $\frac{\delta k^{2}}{\beta}\ll 1$. Second, the lattice has finite size (typically 50x50). This is equivalent to a momentum space spreading of order $0.02$. Since this is within all previous bounds, it should not have a significant affect.
In Section V, we numerically confirm that our techniques are resilient to such physical considerations. We also point out that Ref. [34] experimentally demonstrates that such paths which slightly ‘miss’ the Dirac point, and gap openings due to effects such as the vector a.c. Stark shift, do not significantly affect the physics. Further, they explicitly demonstrate the coherence evolution of the wavevector after it has passed through one node, which is also crucial to our techniques.
.3 III. Details for Method 1: Interferometry